# The Fundamental Parameters of Astrophysical Plasma Turbulence and its Dissipation: Nonrelativistic Limit

### v1

@@ 6.2 Turbulent Damping Mechanisms on the $(\beta_{i},k_{0}\rho_{i})$ Plane
Figure 11: Phase diagram for the kinetic damping mechanisms of weakly collisional plasmas turbulence as a function of isotropic driving wavenumber $k_{0}\rho_{i}$ and ion plasma beta $\beta_{i}$, showing regions of this parameter space where different mechanisms are likely to contribute to the heating of the plasma species: ion Landau damping (iLD) and ion transit time damping (iTTD) (blue); electron Landau damping (eLD) and electron transit time damping (eTTD) (blue); ion cycltron damping (iCD) (black); ion stochastic heating (iSH) (green); ion kinetic viscous heating (iVH) (magenta); and collisionless magnetic reconnection (RXN) (red) for intermittent current sheets ($n=1$, red dashed) and sinusoidal current sheets ($n=2$, red solid). The extent of the turbulent near-Earth solar wind on the $(\beta_{i},k_{0}\rho_{i})$ plane is indicated (gray shading).
A single phase diagram that demonstrates the power of using the dimensionless plasma and turbulence parameters for the isotropic temperature case presented in Table 3 is a plot of the predicted regions on the $(\beta_{i},k_{0}\rho_{i})$ plane where different kinetic mechanisms are expeected to play a role the damping of weakly collisional plasma turbulence, presented here in Fig. 11.
For this phase diagram, we specify the ion-to-electron temperature ratio $\tau=1$ and the ion-to-electron mass ratio $\mu=1836$. For the isotropic temperature case, the species temperature anisotropies $A_{i}=A_{e}=1$, and we take the turbulence to be driven strongly ($\chi_{0}=0$) and isotropically ($k_{\parallel 0}/k_{\perp 0}=1$) so that the turbulent driving can be characterized by the isotropic driving wavenumber $k_{0}\rho_{i}$. Furthermore, we assume the turbulence driving to be balanced ($Z_{0}^{+}/Z_{0}^{-}=1$) and incompressible ($E_{comp}/E_{inc}=0$). In addition, the turbulence is assumed to be weakly collisional with $k_{\parallel 0}\lambda_{mfp,e}\ll 1$, eliminating magnetic pumping as a potential turbulent damping mechanism. Finally, we specifically choose to use the B06 scaling for MHD turbulence with $\alpha=1$ to assess the contributions of the different proposed damping mechanisms.
@@
For the Landau-resonant ($n=0$) collisionless damping mechanisms—Landau damping and transit-time damping—the primary parameter dependencies summarized in §5.4 are $\beta_{i}$ for ion Landau damping (iLD) and ion transit-time damping (iTTD) and $(\beta_{i},\tau,\mu)$ for electron Landau damping (eLD) and electron transit-time damping (eTTD). Since the sum of LD and TTD for each species is always leads to damping for isotropic equilibrium temperatures, but seperately the net energy transfer by either mechanism to that species over a full wave period can be positive or negative—as discussed in §5.4 and illustrated in Fig. 7— we do not separate the contributions LD and TTD for a single species. Furthermore, since any turbulent cascade energy not deposited onto the ions by iLD and iTTD at the ion scales $k_{\perp}\rho_{i}\sim 1$ will cascade to $k_{\perp}\rho_{i}\gg 1$ and ultimately be transferred to the electrons, the partitioning of energy between ions and electrons by these Landau-resonant mechanisms is simply a function of $\beta_{i}$. Solutions of the linear Vlasov-Maxwell dispersion relation for the normalized ion damping rate by iLD and iTTD, $(\gamma_{iLD}+\gamma_{iLD})/\omega$ indicate that ion damping is non-neglible for $\beta_{i}\gtrsim 0.5$, so iLD and iTTD will contribute in this region (blue line with right arrow), independent of the drivign scale $k_{0}\rho_{i}$. For $\beta_{i}\lesssim 0.5$, eLD and eTTD will dominate the Landau-resonant damping (blue line with left arrow), but eLD and eTTD will still contribute to terminate the remaining turbulent cascade energy not removed by ions (blue right arrow).
@@
Kinetic viscous heating of ions (iVH) mediated by temperature anisotropy instabilities a high values of ion plasma beta $\beta_{i}\gg 1$ depends on the $\beta_{i}$ and the amplitude of the turbulent fluctuations at the driving scale, as shown by (42) in §5.8. Because this mechanism only contributes when $\beta_{i}$ is sufficiently high that the temperature anisotropy exceeds the threshold values for the proton temperature anisotropy instabilities, we take this kinetic damping mechanism as dependent only on the value of the ion plasma beta, with an estimated onset at $\beta_{i}\gtrsim 10$ (magenta line and right arrow).
@@
Finally, if the driving scale $k_{0}\rho_{i}\ll 1$ and the ion plasma beta $\beta_{i}\ll 1$ of the turbulence are sufficiently low, current sheets with a large width-to-thickness ratio in the plane perpendicular to the equilibrium magnetic field can be unstable to a rapid collisionless tearing instability in the large-guide-field limit, as detailed in §5.9. If growth rate of the tearing instability is faster than the rate of the nonlinear cascade of energy to small scales by the turbulence $\gamma_{RXN}/\omega{nl}\gtrsim 1$, collisionless magnetic reconnection can disrupt the turbulent cascade. The condition that the normalized collisionless tearing growth rates satisfy $\gamma_{RXN}/\omega\gtrsim 1$ occurs for the maximum growth rate of the tearing instability is given by either (52) or (57) (both expressions yield that same maximum value at the transitional value of $\Delta^{\prime}\delta_{in}=1$). In these expressions, $n=1$ corresponds to a Harris-like current sheet with a hyperbolic tangent profile, and $n=2$ corresponds to a sinusoidal variation of the reconnection magnetic field. The determination of the tearing growth rate demands $k_{\perp}\rho_{i}\ll 1$ for the validity of this linear tearing growth rate calculation, so we take $k_{\perp}\rho_{i}\rightarrow\mathcal{F}$, where $\mathcal{F}\ll 1$ is the maximum value of the normalized perpendicular wavenumber for which the growth rate derviations remain valid. Here, we specify a value $\mathcal{F}=0.3$ for our estimation of the onset of collisionless magnetic reconnection, plotting the boundaries for reconnection in intermittent current sheets ($n=1$) (dashed red line and downward arrow) and for reconnection in sinusoidal current sheets ($n=2$) (solid red line and downward arrow). Note that the dependence of $k_{0}\rho_{i}$ on $\mathcal{F}$ in both (52) and (57) is very strong, scaling as $\mathcal{F}^{9}$ for $n=1$ and as $\mathcal{F}^{7}$ for $n=2$, so the determination of the maximal value of $k_{\perp}\rho_{i}\equiv\mathcal{F}$ for the tearing instability growth has a strong influence on whether reconnection will arise.
@@
The phase diagram for the turbulent damping mechanisms on the $(\beta_{i},k_{0}\rho_{i})$ plane in Fig. 11 can be used in the following manner to predict which damping mechanisms will contribute to the dissipation of turbulence in a specific space or astrophysical plasma system. In the near-Earth solar wind, the observed ion plasma beta values span the range $0.05\lesssim\beta_{i}\lesssim 10$ (Wilson et al., 2018), taking a 5% and 95% limit on the distribution of $\beta_{i}$ values. Furthermore, the isotropic driving wavenumber in the near-Earth solar wind spans the typical values $10^{-4}\lesssim k_{0}\rho_{i}\lesssim 10^{-3}$ (Tu & Marsch, 1990, 1995; Matthaeus et al., 2005; Howes et al., 2008a; Kiyani et al., 2015). This range of parameters for the solar wind at 1 AU is plotted as the gray shaded region on Fig. 11. Reading the phase diagram, the prediction for the damping mechanisms in the near-Earth solar wind is that Landau-resonant damping for ions (iLD and iTTD) and electrons (eLD and eTTD) will contribute to the damping over the full range of solar wind turbulence parameters. At lower values of $\beta_{i}\lesssim 0.3$ and higher values of $k_{0}\rho_{i}$, stochastic heating is predicted to play a role. Indeed, in the Earth’s magnetosheath (which has a similar footprint on the $(\beta_{i},k_{0}\rho_{i})$ plane as the near-Earth solar wind), there is direct evidence from Magnetospheric MultiScale spacecraft observations of electron Landau damping playing a significant role in the damping of the turbulent fluctuations (Chen et al., 2019; Afshari et al., 2021). In addition, in the solar wind in the inner heliosphere, where the values of $\beta_{i}$ tend to be lower than in the near-Earth solar wind, analysis of observations from the Helios spacecraft (Martinović et al., 2019) and Parker Solar Probe spacecraft (Martinović et al., 2020) find evidence for ion stochastic heating in the turbulent plasma. Thus, phase diagrams like Fig. 11 can play a valuable role in predicting the mechanisms that will contribute to the damping of weakly collisional plasma turbulence in space and astrophysical plasmas.
@@ 7 Conclusions
$(\beta_{i},\tau,A_{i},A_{e},k_{\parallel 0}\lambda_{mfp,e};k_{\perp 0}\rho_{i},k_{\parallel 0}/k_{\perp 0},\chi_{0},Z_{0}^{+}/Z_{0}^{-},E_{comp}/E_{inc}),$
summarized in Table 2. Although this set of ten parameters is sufficient to characterize a wide variety of turbulent space and astrophysical plasmas, the development of completely general turbulent heating models on a ten-dimensional parameter space is unlikely to be successful. Fortunately, we can capture many of the dominant dependencies of the turbulence and its damping mechanisms by adopting of reduced set of parameters in the limits that (i) the equilibrium velocity distribution for each species is an isotropic Maxwellian velocity and (ii) the turbulence is driven at sufficiently large scale or sufficiently strongly that the turbulent fluctuations at the small-scale end of the inertial range $k_{\perp}\rho_{i}\sim 1$ satisfy a state of strong turbulence, $\chi\sim 1$. With these idealizations to the isotropic temperature case, there remain just three dimensionless parameters
@@
suumarized in Table 3.
A number of kinetic mechanisms have been proposed to govern the damping of weakly collisional plasma turbulence, as enumerated in §5.1. The critical advance presented in this paper is to express the dependence of each of these mechanisms on the same set of fundamental plasma and turbulence parameters. The results of this analysis of the dependencies for each of the turbulent damping mechanisms is summarized in Table 4.
Finally, the scalings of each of the turbulent damping mechanisms with the isotropic driving wavenumber $k_{0}\rho_{i}$ and the ion plasma beta $\beta_{i}$ can be synthesized to generate the first phase diagram for turbulent damping mechanisms as a function of $(\beta_{i},k_{0}\rho_{i})$, presented in Fig. 11. For a chosen space or astrophysical plasma system, such as the near-Earth solar wind, the range of dimensionless plasma and turbulence parameters can be identified on plotted on the phase diagram, as illustrated by the gray shaded region in the figure. One may then read off which mechnisms are predicted to contribute to the damping of the turbulence in that system. Kinetic numerical simulations and spacecraft observations of weakly collisional turbulence in space or astrophysical plasmas can be used to update and refine this phase diagram, or to extend its applicability as additional dimensionless parameters are varied, such as the ion-to-electron temperature ratio $\tau=T_{i}/T_{e}$.
Ulimately, the goal is to produce more accurate and complete predictive turbulent heating models than those reviewed in §4, facilitating improved global modeling of important turbulent space and astrophysical plasmas, such as the mesoscale and macroscale energy transport through the heliosphere or the interpretation of Event Horizon Telescope observations of accretion disks around supermassive black holes.
Acknowledgements
More than fifteen years of turbulent modeling efforts are synthesized here, having benefitted from interactions over many years with Steve Cowley, Bill Dorland, Eliot Quataert, Alex Schekochihin, Greg Hammett, Tomo Tatsuno, Ryusuke Numata, Ben Chandran, Stuart Bale, Kevin Nielson, Jason TenBarge, Jan Drake, Fred Skiff, Craig Kletzing, Steve Vincena, Troy Carter, Chadi Salem, Kristopher Klein, Chris Chen, Jim Schroeder, Daniel Told, Frank Jenko, Foaud Sahraoui, Tak Chu Li, Suranga Ruhunusiri, Jasper Halekas, Sofiane Bourouaine, Jaye Verniero, Matt Kunz, Andrew McCubbin, Francesco Valentini, Robert Wicks, Daniel Verscharen, Lloyd Woodham, Sarah Conley, Arya Afshari, Peter Montag, Daniel McGinnis, Rui Huang, and Alberto Felix.

### v2

@@ 6.2 Turbulent Damping Mechanisms on the $(\beta_{i},k_{0}\rho_{i})$ Plane
Figure 12: Phase diagram for the kinetic damping mechanisms of weakly collisional plasmas turbulence as a function of isotropic driving wavenumber $k_{0}\rho_{i}$ and ion plasma beta $\beta_{i}$, showing regions of this parameter space where different mechanisms are likely to contribute to the heating of the plasma species: ion Landau damping (iLD) and ion transit time damping (iTTD) (blue); electron Landau damping (eLD) and electron transit time damping (eTTD) (blue); ion cyclotron damping (iCD) (black); ion stochastic heating (iSH) (green); ion kinetic viscous heating (iVH) (magenta); and collisionless magnetic reconnection (RXN) (red) for intermittent current sheets ($n=1$, red dashed) and sinusoidal current sheets ($n=2$, red solid). The extent of the turbulent near-Earth solar wind on the $(\beta_{i},k_{0}\rho_{i})$ plane is indicated (gray shading).
A single phase diagram that demonstrates the power of using the dimensionless plasma and turbulence parameters for the isotropic temperature case presented in Table 4 is a plot of the predicted regions on the $(\beta_{i},k_{0}\rho_{i})$ plane where different kinetic mechanisms are expected to play a role the damping of weakly collisional plasma turbulence, presented here in Fig. 12. This phase diagram of the turbulent damping mechanisms is the key result of this study.
For this phase diagram, we specify the ion-to-electron temperature ratio $\tau=1$ and the ion-to-electron mass ratio $\mu=1836$. For the isotropic temperature case, the species temperature anisotropies are $A_{i}=A_{e}=1$, and we take the turbulence to be driven strongly ($\chi_{0}=0$) and isotropically ($k_{\parallel 0}/k_{\perp 0}=1$) so that the turbulent driving can be characterized by the isotropic driving wavenumber $k_{0}\rho_{i}$. Furthermore, we assume the turbulence driving to be balanced ($Z_{0}^{+}/Z_{0}^{-}=1$) and incompressible ($E_{comp}/E_{inc}=0$). In addition, the turbulence is assumed to be weakly collisional with $k_{\parallel 0}\lambda_{mfp,e}\ll 1$, eliminating magnetic pumping as a potential turbulent damping mechanism. Finally, we specifically choose to use the B06 scaling for MHD turbulence with $\alpha=1$ to assess the contributions of the different proposed damping mechanisms.
@@
For the Landau-resonant ($n=0$) collisionless damping mechanisms—Landau damping and transit-time damping—the primary parameter dependencies summarized in §5.4 are $\beta_{i}$ for ion Landau damping (iLD) and ion transit-time damping (iTTD) and $(\beta_{i},\tau,\mu)$ for electron Landau damping (eLD) and electron transit-time damping (eTTD). Since the sum of LD and TTD for each species always leads to damping in the case of isotropic equilibrium temperatures, but separately the net energy transfer by either mechanism to that species over a full wave period can be positive or negative—as discussed in §5.4 and illustrated in Fig. 7— we do not separate the contributions of LD and TTD for a single species. Furthermore, since any turbulent cascade energy not deposited onto the ions by iLD and iTTD at the ion scales $k_{\perp}\rho_{i}\sim 1$ will cascade to $k_{\perp}\rho_{i}\gg 1$ and ultimately be transferred to the electrons, the partitioning of energy between ions and electrons by these Landau-resonant mechanisms is simply a function of $\beta_{i}$. Solutions of the linear Vlasov-Maxwell dispersion relation for the normalized ion damping rate by iLD and iTTD, $(\gamma_{iLD}+\gamma_{iTTD})/\omega$ indicate that ion damping is non-negligible for $\beta_{i}\gtrsim 0.5$, so iLD and iTTD will contribute in this region (blue line with right arrow), independent of the driving scale $k_{0}\rho_{i}$. For $\beta_{i}\lesssim 0.5$, eLD and eTTD will dominate the Landau-resonant damping (blue line with left arrow), but eLD and eTTD will still contribute to terminate the remaining turbulent cascade energy not removed by ions at $\beta_{i}\gtrsim 0.5$ (blue right arrow).
@@
Kinetic viscous heating of ions (iVH) mediated by temperature anisotropy instabilities [^16] Recall that this mechanism is not the standard viscous heating mediated by microscopic Coulomb collisions between charged plasma particles, as emphasized in §5.8. at high values of ion plasma beta $\beta_{i}\gg 1$ depends on the $\beta_{i}$ and the amplitude of the turbulent fluctuations at the driving scale, as shown by (46) in §5.8. Because this mechanism only contributes when $\beta_{i}$ is sufficiently high that the temperature anisotropy can exceed the threshold values for the proton temperature anisotropy instabilities given by Table 6 and shown in Fig. 8, we take the threshold for this kinetic damping mechanism to be dependent only on the value of the ion plasma beta, with an estimated onset at $\beta_{i}\gtrsim 10$ (magenta line and right arrow).
@@
Finally, if the isotropic driving wavenumber $k_{0}\rho_{i}\ll 1$ and the ion plasma beta $\beta_{i}\ll 1$ of the turbulence are sufficiently low, current sheets with a large width-to-thickness ratio in the plane perpendicular to the equilibrium magnetic field can be unstable to a rapid collisionless tearing instability in the large-guide-field limit, as detailed in §5.9. If growth rate of the tearing instability is faster than the rate of the nonlinear cascade of energy to small scales in the turbulence $\gamma_{RXN}/\omega_{nl}\gtrsim 1$, collisionless magnetic reconnection can disrupt the turbulent cascade. The condition that the normalized collisionless tearing instability growth rate satisfies $\gamma_{RXN}/\omega\gtrsim 1$ (for the unstable wavenumber with the maximum growth rate) is given by either (56) or (61) (both expressions yield the same maximum value at the transitional value of $\Delta^{\prime}\delta_{in}=1$). In these expressions, $n=1$ corresponds to a Harris-like current sheet with a hyperbolic tangent profile, and $n=2$ corresponds to a sinusoidal variation of the reconnecting magnetic field. The determination of the tearing growth rate demands $k_{\perp}\rho_{i}\ll 1$ for the validity of this linear tearing growth rate calculation, so we take $k_{\perp}\rho_{i}\rightarrow\mathcal{F}$, where $\mathcal{F}\ll 1$ is the maximum value of the normalized perpendicular wavenumber for which the growth rate derivations remain valid. Here, we specify a value $\mathcal{F}=0.3$ for our estimation of the onset of collisionless magnetic reconnection, plotting the threshold boundaries for reconnection of intermittent current sheets ($n=1$) (dashed red line and downward arrow) and for reconnection of sinusoidal current sheets ($n=2$) (solid red line and downward arrow). Note that the dependence of $k_{0}\rho_{i}$ on $\mathcal{F}$ in both (56) and (61) is very strong, scaling as $\mathcal{F}^{9}$ for $n=1$ and as $\mathcal{F}^{7}$ for $n=2$, so the determination of the maximal value of $k_{\perp}\rho_{i}\equiv\mathcal{F}$ for the tearing instability growth has a strong influence on whether reconnection will arise.
@@
The phase diagram for the turbulent damping mechanisms on the $(\beta_{i},k_{0}\rho_{i})$ plane in Fig. 12 can be used in the following manner to predict which damping mechanisms will contribute to the dissipation of turbulence in a specific space or astrophysical plasma system. In the near-Earth solar wind, the observed ion plasma beta values span the range $0.05\lesssim\beta_{i}\lesssim 10$ (Wilson et al., 2018), taking a 5% and 95% limit on the distribution of $\beta_{i}$ values. Furthermore, the isotropic driving wavenumber in the near-Earth solar wind spans the typical values $10^{-4}\lesssim k_{0}\rho_{i}\lesssim 10^{-3}$ (Tu & Marsch, 1990, 1995; Matthaeus et al., 2005; Howes et al., 2008a; Kiyani et al., 2015). This range of parameters for the solar wind at 1 AU is plotted as the gray shaded region on Fig. 12. Reading the phase diagram, the prediction for the damping mechanisms in the near-Earth solar wind is that Landau-resonant damping for ions (iLD and iTTD) and electrons (eLD and eTTD) will contribute to the damping over the full range of solar wind turbulence parameters. At lower values of $\beta_{i}\lesssim 0.3$ and higher values of $k_{0}\rho_{i}$, stochastic heating is predicted to play a role. Indeed, in the Earth’s magnetosheath (which has a similar footprint on the $(\beta_{i},k_{0}\rho_{i})$ plane as the near-Earth solar wind), there is direct evidence from Magnetospheric MultiScale spacecraft observations of electron Landau damping playing a significant role in the damping of the turbulent fluctuations (Chen et al., 2019; Afshari et al., 2021). In addition, for the solar wind in the inner heliosphere, where the values of $\beta_{i}$ tend to be lower than in the near-Earth solar wind, analysis of observations from the Helios spacecraft (Martinović et al., 2019) and Parker Solar Probe spacecraft (Martinović et al., 2020) find evidence for ion stochastic heating in the turbulent plasma. Thus, phase diagrams like Fig. 12 can play a valuable role in predicting the mechanisms that will contribute to the damping of weakly collisional plasma turbulence in space and astrophysical plasmas.
@@ 7 Conclusions
$(\beta_{\parallel,i},\tau_{\parallel},A_{i},A_{e},k_{\parallel 0}\lambda_{mfp,e};k_{\perp 0}\rho_{i},k_{\parallel 0}/k_{\perp 0},\chi_{0},Z_{0}^{+}/Z_{0}^{-},E_{comp}/E_{inc}),$
summarized in Table 3. Although this set of ten parameters is sufficient to characterize a wide variety of turbulent space and astrophysical plasmas, the development of completely general turbulent heating models on a ten-dimensional parameter space is unlikely to be successful. Fortunately, we can capture many of the dominant dependencies of the turbulence and its damping mechanisms by adopting of reduced set of parameters in the limits that (i) the equilibrium velocity distribution for each species is an isotropic Maxwellian and (ii) the turbulence is driven at sufficiently large scale or sufficiently strongly that the turbulent fluctuations at the small-scale end of the inertial range $k_{\perp}\rho_{i}\sim 1$ satisfy a state of strong turbulence, with a nonlinearity parameter at those small scales of $\chi\sim 1$. With these idealizations we arrive at the isotropic temperature case, where just three dimensionless parameters remain
@@
summarized in Table 4.
A number of kinetic mechanisms have been proposed to govern the damping of weakly collisional plasma turbulence, as enumerated in §5.1. The critical advance presented in this paper is to express the dependence of each of these mechanisms on the same set of fundamental plasma and turbulence parameters. The result of this analysis of the dependencies for each of the turbulent damping mechanisms is summarized in Table 5.
Finally, the scalings of each of the turbulent damping mechanisms with the isotropic driving wavenumber $k_{0}\rho_{i}$ and the ion plasma beta $\beta_{i}$ can be synthesized to generate the first phase diagram for turbulent damping mechanisms as a function of $(\beta_{i},k_{0}\rho_{i})$, presented in Fig. 12. For a chosen space or astrophysical plasma system, such as the near-Earth solar wind, the range of dimensionless plasma and turbulence parameters can be plotted on the phase diagram, as illustrated by the gray shaded region in the figure. One may then read off which mechanisms are predicted to contribute to the damping of the turbulence in that system. Kinetic numerical simulations and spacecraft observations of weakly collisional turbulence in space or astrophysical plasmas can be used to update and refine this phase diagram, or to extend its applicability as additional dimensionless parameters are varied, such as the ion-to-electron temperature ratio $\tau=T_{i}/T_{e}$. It is worthwhile emphasizing if kinetic simulations are used to identify the turbulent damping mechanisms and to quantify the associated energy density transfer rates to each plasma species, it is essential that those simulations be performed in the full three spatial dimensions. Utilizing reduced dimensionality for such simulations may unphysically restrict the available channels of energy transfer from the turbulent fluctuations to the particles, potentially changing which physical turbulent damping mechanisms are dominant and thereby artificially altering the resulting partitioning of energy among particle species and degrees of freedom.
Ultimately, the goal is to produce more accurate and complete predictive turbulent heating models than those reviewed in §4, facilitating improved global modeling of important turbulent space and astrophysical plasmas, such as the mesoscale and macroscale energy transport through the heliosphere or the interpretation of Event Horizon Telescope observations of accretion disks around supermassive black holes.
Acknowledgments
More than fifteen years of turbulent modeling efforts are synthesized here, having benefited from interactions over many years with Steve Cowley, Bill Dorland, Eliot Quataert, Alex Schekochihin, Greg Hammett, Tomo Tatsuno, Ryusuke Numata, Ben Chandran, Stuart Bale, Kevin Nielson, Jason TenBarge, Jan Drake, Fred Skiff, Craig Kletzing, Steve Vincena, Troy Carter, Chadi Salem, Kristopher Klein, Chris Chen, Jim Schroeder, Daniel Told, Frank Jenko, Foaud Sahraoui, Tak Chu Li, Suranga Ruhunusiri, Jasper Halekas, Sofiane Bourouaine, Jaye Verniero, Matt Kunz, Andrew McCubbin, Francesco Valentini, Robert Wicks, Daniel Verscharen, Lloyd Woodham, Sarah Conley, Arya Afshari, Peter Montag, Daniel McGinnis, Collin Brown, Rui Huang, and Alberto Felix.